up date

Mahnaz mahnaz at snf.stanford.edu
Wed Jul 24 16:45:18 PDT 2002


hello all,

Our Afm head came back  and yesterday Juan fiddle with it and no success
and today after Hector looking at it we realized that the laser
adjustments screw are not functioning correctly. Thanks to Tony we
shipped the part back this afternoon and DI will send out a loaner
tomorrow.

mahnaz




More information about the afm2 mailing list